# SteerMusic: Enhanced Musical Consistency for Zero-shot Text-Guided and Personalized Music Editing

### v1

@@ 5.4. Personalized Music Editing
Baseline: We set two existing personalized music editing methods proposed by (Plitsis et al., 2024) as the baselines, Textual inversion and DreamSound. Textual inversion optimizes a concept embedding, whereas DreamSound fine-tunes a AudioLDM2 with rare-token identifiers (Ruiz et al., 2023). Both methods perform personalized music editing by manipulating the concept token during denoising process. We follow the official codes provided by (Plitsis et al., 2024) to obtain text-to-music PDM1. According to (Plitsis et al., 2024), we reproduce the personalized music editing methods by calculating noisy latent representations $x_{t}$ from $x_{0}^{\text{src}}$ of a DPM conditioned on the source prompt with a predefined shallow time step $t$ using DDIM inversion (Song et al., [n. d.]), where $t=30$. We denoise $x_{t}$ on the PDM [^5] Since textual inversion optimizes only the concept token embedding rather than fine-tuning a PDM, we denoise $x_{t}$ using the DPM employed during DDIM inversion. Consequently, textual inversion is not compatible with SteerMusic+. conditioned on the target prompt linked to the learned concept to obtain $x_{0}^{\text{tgt}}$. We do not compare with Jen-1 DreamStyler (Chen et al., 2024), as it focuses on personalized music generation rather than editing.
@@
Experimental results: We plug-in SteerMusic+ in the same PDM used in DreamSound5. Table 2 shows that SteerMusic+ achieves superior musical consistency compare to the two baselines. It indicates that the edited outputs of SteerMusic+ successfully preserve instruction-irrelevant music content in the source music. Furthermore, SteerMusic+ performs accurate edits that align well with the concepts captured from references as indicated by the low CDPAM compared to the baseline methods. These objective results are aligned to the subjective metrics, where SteerMusic+ obtains higher MOS-P and MOS-T than the baselines. Figure 8 provides a visual comparison on personalized music instrument transfer task. SteerMusic+ successfully preserves instruction-irrelevant music content on the source music, such as vocal information. More visualizations are available in our supplementary material.
@@
Figure 9. User preference study for SteerMusic+. We report the percentage of users that preferred SteerMusic+ (ours) over DreamSound and Textual inv. (Plitsis et al., 2024) in this user study.
@@
User preference study: Similar to Sec. 5.3, we evaluate SteerMusic+ by a user preference study compared to DreamSound and Textual inversion2. In this study, users were asked to select the edit result that the style better matches the reference music style while preserving the main source music content. We collect 24 full responses, which participants have minimum of 1 and on average 5 years of music training experience. As shown in Figure 9, SteerMusic+ are clearly preferred by users than the baselines.
@@
Figure 10. Ablation study on SteerMusic+ with adherence to music consistency vs. edit fidelity on the edited music with different $\lambda$ values in Eq. 8. The horizontal axis (CQT-1 PCC) measures how well the source melody is preserved, while the vertical axis (CDPAM) reflects closeness to the target concept.
@@
Ablation study: We conduct an ablation study to understand the effects of different components in SteerMusic+. We used the concept [bouzouki] as it is an uncommon instrument style that typically requires personalized models. Yellow dots indicate results obtained by $\mathcal{L}_{\text{PDS-O}}$ in Eq. 8 alone, with varying $\lambda$ values. From the zoomed-in view in Figure 10, we observe that as the CDPAM value decreases with $\lambda$, the loss of $\mathcal{L}_{\text{shift}}$ within $\mathcal{L}_{PDS-O}$ helps steer the editing process toward the personalized concept. However, the effect of $\mathcal{L}_{\text{shift}}$ is minor, suggesting that the gradient of $\mathcal{L}_{\text{PDS}}$ alone does not sufficiently capture the target concept during editing. As a result, the edited music struggles to align with the reference musical characteristics. In contrast, SteerMusic+ (with $\mathcal{L}_{\text{PCon}}$ in Eq. 9) leads to a significant decrease in CDPAM value, indicating better alignment with the target concept. Since $\mathcal{L}_{\text{PCon}}$ was proposed to explicitly enhance the characteristics of the target concept on the editing, we observe that $\mathcal{L}_{\text{shift}}$ becomes more effective in this setting. Specifically, when $\lambda$ is negative, the edit preserves the original musical content, while positive $\lambda$ push the edit toward the target concept.
@@
This reveals a fundamental trade-off between music consistency and adherence to the target style. In personalized music editing, where musical characteristics are derived from reference music, shifting the edit closer to the concept inevitably disrupts the original content. This is distinct from text-guided music editing, where the transformation follows a textual description rather than an explicit reference. In personalized editing, the challenge lies in balancing the retention of the original musical structure, while integrating the distinctive attributes of the reference music. We suggest that the $\lambda$ value should not be too large for a further regularization in real applications (e.g., $\lambda=0.05$).
@@
Figure 11. Adherence to music consistency vs. edit fidelity to edited results with different fine-tune steps in PDM.
Limitation. As in shown Figure 11, we observe that the number of fine-tune steps with a fixed learning rate in the PDM significantly influences the ability of DreamSound and SteerMusic+ to capture the concept of reference music while preserving the original melody. When the PDM uses few fine-tuning steps (e.g., 50), it struggles to fully internalize the reference concept, resulting in higher CDPAM scores and underwhelming edit consistency. Conversely, with an excessively large number of steps (e.g., 200), the PDM tends to overfit, leading to lower CQT-1 PCC and a loss of source music structure. These findings highlight the importance of balancing fine-tuning steps in the PDM to achieve both high edit fidelity and source content preservation on personalized music editing setting.
6. Conclusion
We present two music editing methods, SteerMusic and SteerMusic+, from coarse-grained to fine-grained music editing. To the best of our knowledge, this is the first attempt to fully explore DDS and score distillation technique in the music editing framework. Our methods overcome the limitation of existing methods that lack musical consistency, causing the edited results to differ from the original music composition. In additional to better musical consistency, our methods also produce high-fidelity edited results that align with the target text prompts. We also extended our method beyond text-guided method and proposed a novel pipeline for personalized music editing. This method allows for fine-grained music editing, which extracts the user-defined music style concept token from reference music, producing more precise and accurate editing. We systematically evaluate our method through quantitative and qualitative metrics, complemented by comprehensive ablation studies. In our experiment, we demonstrated that our methods outperform existing baseline methods, which improves the musical consistency while maintaining high editing fidelity.

### v3

@@ Introduction
Baseline. We set two existing personalized music editing methods proposed by plitsis2024investigating as the baselines, Textual inversion and DreamSound. Textual inversion optimizes a concept embedding, whereas DreamSound fine-tunes an AudioLDM2 with rare-token identifiers (ruiz2023dreambooth). Both methods perform personalized music editing by manipulating the concept token during denoising process. We follow the official codes provided by plitsis2024investigating to obtain text-to-music PDM. We reproduce the personalized music editing methods by calculating noisy latent representations $x_{t}$ from $x_{0}^{\text{src}}$ of a DPM conditioned on the source prompt with a predefined shallow time step $t$ using DDIM inversion (song2020denoising), where $t=30$. We denoise $x_{t}$ on the PDM  [^1] Textual inversion optimizes only the concept token embedding rather than fine-tuning a PDM, we denoise $x_{t}$ using the DPM employed during DDIM inversion. Consequently, textual inversion is incompatible with SteerMusic+ pipeline, which relies on a PDM. conditioned on the target prompt linked to the learned concept to obtain $x_{0}^{\text{tgt}}$. To ensure statistical reliability, experiments are conducted using multiple random seeds. We do not compare with Jen-1 DreamStyler (chen2024jen), which is a personalized music generation method.
@@
Experimental results. We plug SteerMusic+ into the PDM used in DreamSound1. Tab. 2 shows that SteerMusic+ achieves superior musical consistency compared to the baselines. It indicates that the edited outputs of SteerMusic+ successfully preserve instruction-irrelevant music content in the source music. Furthermore, SteerMusic+ performs accurate edits that align well with the concepts captured from references, as indicated by the low CDPAM compared to the baseline methods. These objective evaluation results aligned with the subjective metrics, with SteerMusic+ obtaining significantly higher MOS-P and MOS-T than the baselines as measured by an ANOVA test (MOS-P with p-value $1.38\times 10^{-13}$ and MOS-T with p-value $1\times 10^{-3}$). In long-form music editing (Tab. 3), SteerMusic+ still outperforms the baseline. We exclude Textual Inv. from comparison as it fails to produce meaningful results on MusicDelta, likely due to the limitation of its base model with complex inputs. Fig. 6 shows a personalized instrument style transfer example, where SteerMusic+ effectively preserves instruction-irrelevant content from the source music.
@@
Figure 7: User preference for SteerMusic+: percentage of users preferring our method over the baselines, DreamSound and Textual Inv.
@@
User preference study. We evaluate SteerMusic+ by a user preference study compared to DreamSound and Textual Inv. Users were asked to select the edited result that best matches the reference style while preserving the main content of the source music. We collected 24 full responses, excluding responses that are partially finished. Participants have a minimum of 1 and an average of 5 years of music training experience. As shown in Fig. 7, SteerMusic+ is clearly preferred by participants compared to baselines.
@@
Figure 8: Ablation study on SteerMusic+ with adherence to music consistency vs. edit fidelity on the edited music with $\lambda$ values in Eq. 4. The horizontal axis (CQT1-PCC) indicates source melody preservation; the vertical axis (CDPAM) indicates alignment with the target concept.
@@
Ablation study. We conduct an ablation study to understand the effects of different components in SteerMusic+. We used the concept [bouzouki] as it is an uncommon musical instrument that typically requires the use of personalized models. The regularization weight $\lambda$, which acts as the Lagrange multiplier for the constraint in Eq. 4, was tested within [-1,1] to balance the constraint enforcement and the optimization stability. Yellow dots indicate results obtained by $\mathcal{L}_{\text{PDS-O}}$ in Eq. 4 with varying $\lambda$ values. From the zoomed-in view in Fig. 8, as the CDPAM value decreases with $\lambda$, the loss of $\mathcal{L}_{\text{shift}}$ within $\mathcal{L}_{\text{PDS-O}}$ helps steer the editing toward the concept. However, the effect of $\mathcal{L}_{\text{shift}}$ is minor, suggesting that the gradient of $\mathcal{L}_{\text{PDS}}$ alone does not sufficiently capture the target concept during editing. As a result, the edited music struggles to align with the reference musical characteristics. In contrast, SteerMusic+ (with $\mathcal{L}_{\text{PCon}}$ in Eq. 5) leads to a significant decrease in the CDPAM value, which indicates better alignment with the target concept. Since $\mathcal{L}_{\text{PCon}}$ was proposed to explicitly enhance the characteristics of the target concept on the editing, we observe that $\mathcal{L}_{\text{shift}}$ becomes more effective in this setting. Specifically, when $\lambda$ is negative, the edit preserves the original content, while positive $\lambda$ pushes the edit toward the target concept.
@@
This highlights a fundamental trade-off between the source music consistency and the adherence to the target style, consistent with the findings of manor2024zero. In personalized music editing, where target characteristics are derived from reference music, shifting the output towards the concept often disrupts the original content. This differs from text-guided editing, which follows abstract textual cues rather than concrete musical references. The key challenge in personalized editing is balancing the retention of the original music content, while integrating the distinctive attributes of the reference music. For practical use, we recommend keeping a smaller $\lambda$ value (e.g., $\lambda=0.05$) to avoid over-editing.
@@
Figure 9: Adherence to music consistency vs. edit fidelity to edited results with different fine-tune steps in PDM. The horizontal axis (CQT1-PCC) indicates source melody preservation; the vertical axis (CDPAM) indicates alignment with the target concept.
Limitation. As shown in Fig. 9, the number of fine-tuning steps in the PDM significantly affects the performance of DreamSound and SteerMusic+. A small number of steps (e.g., 50) leads to poor concept learning and low edit fidelity, while a large number of steps (e.g., 200) causes overfitting and source music structural loss. These findings highlight the importance of balancing PDM fine-tuning steps to achieve both high edit fidelity and content preservation in personalized music editing settings. The editing fidelity of SteerMusic+ is fundamentally limited by the ability of the PDMs to capture the concept attributes. Consequently, failure cases can often be attributed to inherent shortcomings in the PDM’s representational capacity. We refer readers to plitsis2024investigating, which investigated the capacity of PDMs to capture musical attributes and offers insights into improving this capacity.
Conclusion
We present two music editing methods, SteerMusic and SteerMusic+, from coarse-grained to fine-grained music editing. To the best of our knowledge, this is the first work to fully leverage DDS and score distillation in the music editing framework. Our methods address the limitations of prior approaches by enhancing musical consistency while producing high-fidelity edits aligned with target prompts. SteerMusic+ further introduces a personalized editing pipeline that extracts user-defined style concepts from reference music for fine-grained control. We validate our methods through comprehensive experiments that show clear improvements over existing baselines in both consistency and fidelity. For practical application, the proposed methods could be extended by using a better-trained TTA diffusion backbone with a higher sampling rate to achieve high-fidelity editing results. In future studies, personalized music editing methods could focus on improving user-controllable editing outcomes for more nuanced and expressive edits.